A nursery receives a shipment of 246 trees. The shipment includes pine trees, fruit trees, and maple trees. There are twice as many pine trees as there are fruit trees. The maple trees make up half of the shipment. How many of each type of tree was in this shipment?
3x3 equation.

Answer by rothauserc(4718)

let p be the number of pine trees, f be the number of fruit trees and m be the number of maple trees
:
p = 2f
:
m = 246/2 = 123
:
2f +f +123 = 246
:
2f +f = 123
:
3f = 123
:
f = 41
:
p = 41 * 2 = 82
:
********************************************************************
there are 82 pine trees, 41 fruit trees, and 123 maple trees
********************************************************************
:
check the answer
:
82 +41 +123 = 246
:
123 +123 = 246
:
246 = 246
:
answer checks
: